# Enhancing System-Level Safety in Mixed-Autonomy Platoon via Safe Reinforcement Learning

### 20 Jan 2024 v1

@@ V-B Testing Results
In the testing phase, two safety-critical scenarios that could cause safety-critical failures for the LCC model are simulated. Scenario 1 considers the scenario where the preceding HDV may brake urgently to avoid collisions with some unexpected cutting-in vehicle or crossing pedestrians, which may cause the distance to the CAV behind to be less than the safety distance. Scenario 2 describes a situation where the HDVs following the CAV in the platoon may suddenly accelerate due to human mistakes caused by driving fatigue. Such an instantaneous acceleration could endanger the vehicles ahead of it.
First, we perform the safety-guaranteed region analysis to quantitatively discuss the enhancement of the system-level safety performance for our proposed method. Second, we perform the case study to analyze the safety performance in specific conditions for the two safety-critical scenarios. The setting for the case study is given as follows. For scenario 1, the disturbance signal is added to the preceding HDV for simulation. Consider an emergency happens at 00s that leads to the preceding HDV $(\rm index=0\quad or\quad index=1)$ decelerating with $\rm-4m/s^{2}$ for $2.5$s, and the head HDV maintains the low velocity for another $2.5$s, then speeds up to equilibrium speed after the risk is avoided. In scenario 2, the follower HDVs after the CAV in the platoon may suddenly accelerate due to human mistakes caused by driving fatigue. Such an instantaneous acceleration could endanger the vehicles ahead of it. In particular, we add disturbances to the $3$-rd or $4$-th HDV which is behind the CAV. At 00s, the following HDV accelerates with $\rm 1m/s^{2}$ for $4$s, and maintains high velocity for $4s$, then decelerates to an equilibrium state. Note that this structured and extreme testing scenario is significantly different from the training scenarios with random disturbances. Such treatment is to demonstrate the performance of the safety layer under adversarial conditions and the generalizability of our model to unseen scenarios. We compare four types of models: pure HDVs, PPO, PPO with the safety layer without learning-based human driver behavior identification and PPO with the safety layer and learning-based human driver behavior identification.
@@ V-B1 Safety-Guaranteed Region Analysis
Figure 5: Safety-guaranteed regions associated with two specific scenarios: (a) the deceleration disturbance of the preceding vehicle and (b) the acceleration disturbance of the following vehicle. $x$-axis denotes the value of the disturbance signal (acceleration/deceleration), and the $y$-axis is the duration time of the disturbance signal. The dark blue regions denote the safety region of the LCC when employing the PPO controller, and the light blue regions illustrate the expanded safety region achieved by implementing the safe RL controller for LCC, while the white regions indicate the unsafe region.
To demonstrate the extent to which the safety of the platoon is improved, we draw the safety region for the two safety-critical scenarios. In Fig. 5 (a), the HDV indexed in $1$ decelerates for a range of duration times and deceleration rates. It can be seen that the safety region is expanded by almost 60% using the proposed safe RL controller. Moreover, in Fig. 5 (b), the HDV with index $3$ accelerates within a range of duration times and acceleration rates. Utilizing the proposed method, the following HDV’s safety duration time increases by an average of 0.8s. In summary, our method effectively enhances both safety regions, illustrating a marked improvement in safety for both scenario 1 and scenario 2.
@@ V-B2 Case Study for Scenario 1
Figure 6: Headway for CAV when the preceding vehicle emergency decelerates.
@@
Figure 7: Spacing, velocity and acceleration for vehicles in scenario 1. (a), (b), (c) are the results for PPO with the safety layer. (d), (e), (f) correspond to the results for PPO. The results for the pure HDV scenario are presented in (g), (h), (i).
Fig. 6 displays the headway and Fig. 7 shows the spacing, velocity and acceleration at each time step in the test simulation. In this context, the dynamics of HDVs are excluded from the CAV safety constraint, which means that the online human driver behavior identification module for HDVs does not exert any influence on the outcomes. Consequently, the compared models involve pure HDVs, PPO, and PPO with the safety layer. For the pure HDV setting and PPO with the safety layer setting, the spacing and headway between CAV and the preceding vehicle is larger than 00, which indicates that the collision does not occur. However, for the approach without safety layers, safety is not guaranteed for PPO, since the CAV under the algorithm collides with the leading vehicle. In contrast, the PPO with safety layer ensures collision-free operation throughout the entire testing process. This is achieved through the safety constraint for the CAV, which guarantees that the headway between the CAV and the preceding vehicle remains within the designated safe set. Furthermore, as illustrated in Fig. 7 (c), the control input remains within the actuator limitation bound (depicted by the red dotted line), which is ensured by the adherence to feasibility constraint in Eq. (25).
@@ V-B3 Case Study for Scenario 2
Figure 8: Headway for vehicle $4$ when it accelerates emergently.
Figure 9: Headway for vehicle $5$ when it accelerates emergently.
@@
Figure 10: Spacing, velocity and acceleration for vehicles in scenario 2 when vehicle 4 accelerates. (a), (b), (c) are the results for PPO with the safety layer and learning-based human driver behavior identification. (d), (e), (f) show the results for PPO with the safety layer without online system identification. (g), (h), (i) correspond to the results for PPO. The results for the pure HDV scenario are presented in (j), (k), (l).
@@
Figure 11: Spacing, velocity and acceleration for vehicles in scenario 2 when vehicle 5 accelerates. (a), (b), (c) are the results for PPO with the safety layer and learning-based human driver behavior identification. (d), (e), (f) show the results for PPO with the safety layer without learning-based human driver behavior identification. (g), (h), (i) correspond to the results for PPO. The results for the pure HDV scenario are presented in (j), (k), (l).
The results for the HDV safety scenario are presented in Fig. 8, Fig. 9, Fig. 10 and Fig. 11. In Fig. 10, when vehicle $4$ accelerates, both the pure HDVs platoon and the PPO without a safety layer fail to prevent a collision when the following vehicle (indexed $4$) approaches the preceding vehicle. However, collisions are successfully averted by the approaches incorporating a safety layer. In this instance, the impact of disturbances from inaccurate dynamics on safety is minimal. Given that the safety constraints for the following vehicles are treated as soft constraints with relaxation coefficients in optimization formulation (26), it should be noted that the assurance of the following vehicles’ headway within the safe set is not guaranteed but rather enhanced, as shown in Fig. 8.
In Fig. 11, when vehicle $5$ accelerates, collisions occur in the pure HDVs setting, PPO without the safety layer, and PPO with the safety layer but without online system identification. Notably, the HDV safety is ensured for the PPO with a safety layer and online learning-based human driver behavior identification, showcasing the effectiveness of the proposed method in providing a more accurate estimation of system dynamics for enhanced safety. Similarly, the headway for the following vehicles is enhanced by the proposed method in Fig. 9.
@@ VI Conclusion and Future Work
In this paper, we propose a safe RL-based controller for the mixed-autonomy platoon by integrating the DRL method with a differentiable CBF layer to achieve system-level driving safety. A learning-based human driver behavior identification is utilized to derive the unknown dynamics of surrounding HDVs in mixed traffic scenarios. Traffic efficiency and string stability are taken into consideration by constructing comprehensive reward functions. Results show that our proposed method effectively enhances traffic efficiency and stability while providing system-level safety guarantees in a mixed platoon. Furthermore, the safety layer contributes to expediting training by confining the range of exploration, which can be beneficial to the development and deployment of reinforcement learning algorithms.
This research opens several promising directions for future work. First, considering the number of surrounding vehicles for the CAV is dynamic in real-world traffic, a decision transformer can be integrated into the RL controller to adapt to the dynamic traffic scenario. Second, in scenarios where multiple CAVs operate within a mixed-autonomy platoon, multi-agent RL can be employed to collaboratively improve traffic flow, string stability, and system-level safety. Third, the dynamic and unpredictable nature of real traffic conditions suggests the utility of meta-RL to enhance the robustness of the algorithm, ensuring its effectiveness even in some extreme traffic scenarios.

### 01 Mar 2024 v2

@@ V-B Testing Results
In the testing phase, two safety-critical scenarios that could cause safety-critical failures for the LCC model are simulated. Scenario 1 considers the scenario where the preceding HDV may brake urgently to avoid collisions with some unexpected cutting-in vehicle or crossing pedestrians, which may cause the distance to the CAV behind to be less than the safety distance. Scenario 2 describes a situation where the HDVs following the CAV in the platoon may suddenly accelerate due to human mistakes caused by driving fatigue. Such an instantaneous acceleration could endanger the vehicles ahead of it. Note that both scenarios are significantly different from the training scenarios with random disturbances. Such a treatment is to demonstrate the performance of the safety layer under adversarial conditions and the generalizability of our model to unseen scenarios. For each of these two scenarios, we perform (i) a safety-guaranteed region analysis to quantitatively analyze the system-level safety improvement resulting from our proposed method, considering various acceleration rates and acceleration durations, and (ii) a case study to analyze the safety performance under a specific acceleration rate and acceleration duration for the two safety-critical scenarios.
@@ V-B1 Safety-Guaranteed Region Analysis
Figure 5: Safety-guaranteed regions associated with two specific scenarios: (a) the deceleration disturbance of the preceding vehicle and (b) the acceleration disturbance of the following vehicle. The horizontal axis denotes the magnitude of the disturbance signal (acceleration/deceleration), and the vertical axis is the duration of the disturbance signal. The dark blue regions denote the safety region of the LCC with PPO controller without the safety layer, and the light blue regions illustrate the expanded safety region achieved by implementing the safe RL controller for LCC (i.e., with safety layer), while the white regions indicate the unsafe region.
To demonstrate the safety improvement of the platoon due to the incorporation of the safety layer, we show the resulting safety region with and without the safety layer in both safety-critical scenarios. In Fig. 5 (a), the HDV indexed in $1$ decelerates for a range of duration times and deceleration rates. It can be seen that the safety region is expanded by almost 60% using the proposed safe RL controller. Moreover, in Fig. 5 (b), the HDV with index $3$ accelerates within a range of duration times and acceleration rates. Utilizing the proposed method, the following HDV’s safety duration time increases by an average of 0.8s. In summary, our method effectively improves the safety regions in both scenarios.
@@ V-B2 Case Study for Scenario 1
Figure 6: Value of the CBF candidate for CAV when the preceding vehicle emergency decelerates.
@@
Figure 7: Spacing, velocity and acceleration for vehicles in scenario 1. (a), (b), (c) are the results for safe-RL with SI. (d), (e), (f) correspond to the results for PPO without safety guarantee. The results for the pure car-following scenario are presented in (g), (h), (i).
We simulate an emergency that occurs at 00 s, such that a preceding HDV (with index 0 or 1) decelerates with $\rm-4m/s^{2}$ for $2.5$s, maintains a low velocity for another $2.5$s, and then accelerates to equilibrium speed. Note that since the perturbation only occurs on preceding vehicles that cannot be influenced by the CAV, this scenario does not require considering HDV safety, and hence the online human driver behavior identification module does not exert any influence on the outcomes. Consequently, we only compare three models: (i) pure car following, (ii) PPO without safety guarantee, and (iii) safe-RL with SI. Fig. 6 shows the value of the CBF candidate, and Fig. 7 shows the spacing, velocity, and acceleration at each time step in the test simulation. For pure car following and safe-RL with SI, the resulting spacing between the CAV and the preceding HDV and the value of the CBF candidate are both greater than 00, which indicates that the collision does not occur. However, safety is not guaranteed for PPO without safety guarantee, since the CAV under the algorithm collides with the preceding HDV. Furthermore, as illustrated in Fig. 7 (c), the resulting control input from safe-RL with SI remains within the actuator limitation bound (depicted by the red dotted line), which is ensured by the adherence to feasibility constraint in Eq. (22).
@@ V-B3 Case Study for Scenario 2
Figure 8: Value of the CBF candidate for vehicle $3$ when it accelerates emergently.
Figure 9: Value of the CBF candidate for vehicle $4$ when it accelerates emergently.
@@
Figure 10: Spacing, velocity and acceleration for vehicles in scenario 2 when vehicle 4 accelerates. (a), (b), (c) are the results for safe-RL with SI. (d), (e), (f) show the results for safe-RL without SI. (g), (h), (i) correspond to the results for PPO without safety guarantee. The results for the pure car-following scenario are presented in (j), (k), (l).
@@
Figure 11: Spacing, velocity and acceleration for vehicles in scenario 2 when vehicle 5 accelerates. (a), (b), (c) are the results for safe-RL with SI. (d), (e), (f) show the results for safe-RL without SI. (g), (h), (i) correspond to the results for PPO without safety guarantee. The results for the pure car-following scenario are presented in (j), (k), (l).
We add disturbances to the $3$-rd or $4$-th HDV in the platoon, which is behind the CAV. At time 00 s, the following HDV accelerates with $\rm 1m/s^{2}$ for $4$s, maintains high velocity for $4s$, and then decelerates to an equilibrium state. The results are presented in Fig. 8 – Fig. 11.
In Fig. 10, both pure car following and PPO without safety guarantee fail to prevent a collision when vehicle 3 accelerates to approach the preceding vehicle. However, collisions are successfully avoided by safe-RL with SI and safe-RL without SI. In this case, the safety impact of disturbances resulting from inaccurate dynamics is minimal. Given that the safety constraints for the following HDVs are treated as soft constraints with relaxation coefficients in the optimization formulation (23), it should be noted that the assurance for the value of the CBF candidate for the following HDVs within the safe set is not guaranteed but rather enhanced, as shown in Fig. 8.
In Fig. 11, if vehicle $4$ accelerates, collisions occur when applying pure car following, PPO without safety guarantee, or safe-RL without SI. In contrast, safe-RL with SI can ensure safety, showcasing the importance of a more accurate estimation of the system dynamics to improve safety. Similarly, the value of the CBF candidate for the following vehicles is enhanced by safe-RL with SI, as shown in Fig. 9.
@@ VI Conclusion and Future Work
In this paper, we propose a safe RL-based controller for the mixed-autonomy platoon by integrating the DRL method with a differentiable CBF layer to achieve system-level driving safety. A learning-based human driver behavior identification is utilized to derive the unknown dynamics of surrounding HDVs in mixed traffic scenarios. Traffic efficiency and string stability are taken into consideration by constructing comprehensive reward functions. The simulation results show that our proposed method effectively enhances traffic efficiency and stability while providing system-level safety guarantees in a mixed platoon. Furthermore, the safety layer contributes to expediting training by confining the range of exploration, which can be beneficial to the development and deployment of reinforcement learning algorithms.
This research opens several promising directions for future work. First, we would like to apply a decision transformer to adapt to dynamic traffic scenarios where the number of surrounding vehicles of the CAV is time-varying. Second, it would be interesting to leverage multi-agent RL to extend the proposed framework to scenarios where multiple CAVs collaborate within a mixed-autonomy platoon to improve traffic flow, string stability, and system-level safety. Third, we would like to employ meta-RL to improve the generalizability of the proposed safe-RL algorithm to various unseen environments (e.g., road conditions, weather conditions, preferences over multiple control objectives, etc.). Fourth, the proposed methodology has the potential to be extended to consider other types of roads, such as (i) complex scenarios like merges, diverges, and weaving sections that require accounting for lane changes and (ii) urban roads that require coordination between platoon control and signal control.